OpenStudy (anonymous):

• What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms

OpenStudy (anonymous):

8, -16, 32, -64, 128, -256, 512, -1024, 2048, -4096, 8192, -16384, 32768. -65536, 131072. 87384 is the sum.

OpenStudy (mertsj):

\[Sum _{n}=\frac{a _{1}(1-r ^{n})}{1-r}\]

OpenStudy (mertsj):

\[=\frac{8(1-(-2)^{15}}{1-(-2)}\]

OpenStudy (mertsj):

87384

OpenStudy (anonymous):

8(3)^15/3 = 24^15/3?

OpenStudy (mertsj):

No. Remember the rule...powers first

OpenStudy (mertsj):

\[\frac{8(1-(-32768))}{3}\]

OpenStudy (mertsj):

\[\frac{8(32769)}{3}\]

OpenStudy (mertsj):

\[\frac{262152}{3}\]

OpenStudy (mertsj):

\[87384\]

OpenStudy (anonymous):

ooh thank u so much

OpenStudy (mertsj):

yw